Nadia leaves San Francisco at noon. Deena leaves San Francisco at 12:30pm. Deena is traveling 16 mph faster than Nadia. If Deena passes Nadia at 2:00pm, how fast are they each going?
Answer by lwsshak3(11628)

Nadia leaves San Francisco at noon. Deena leaves San Francisco at 12:30pm. Deena is traveling 16 mph faster than Nadia. If Deena passes Nadia at 2:00pm, how fast are they each going?
***
let x=Nadia's speed
Nadia's travel time=12 noon to 2 pm=2 hours
x+16=Deena's speed
Deena's travel time=12:30 noon to 2 pm=1.5 hours
distance=travel time*speed (same for both)
..
2x=1.5(x+16)
2x=1.5x+24
.5x=24
x=48
x+16=64
how fast are they each going?
Nadia's speed=48 mph
Deena's speed=64 mph